Two resistors with resistance values of 4.5 Ω and 2.3 Ω are connected in series or parallel

across a 30 V potential difference to a light bulb.

a. Calculate the current delivered through the light bulb in the two cases.

b. Draw the circuit connection that will achieve the brightest light bulb.

2. Explain the term internal resistance.

3. State how you agree or disagree with the following statement.

A good circuit cannot have internal resistance


1
Expert's answer
2021-07-23T14:49:01-0400

a.) Current delivered

Series connections

Total resistance for series connections = 4.5+2.3 =6.8 Ω

"I=\frac{V}{R}"

"I=\frac{30}{6.8}=4.4A"

Parallel connections

Total resistance for parallel connections =

"\frac {1}{R_T}=\frac{1}{R_1}+ \frac{1}{R_2}"

"\frac {1}{R_T}=\frac{1}{4.5}+\frac {1}{2.3}=1.5Ω"

"I=\frac{30}{1.5}=20A"


b.) Circuit diagram




c.) Internal resistance is the opposition of the flow of current offered by the cells and batteries resulting in the generation of heat. 


d.) The circuits for appliances such as automatic lights need a lot of internal resistance for the Light Dependant Resistors (LDR) to efficiently work while circuits that are used carrying charge to be used by other appliances need as low internal resistance as possible for it to be efficient.

Therefore I disagrees with this statement that a good circuit cannot have internal resistance